OpenStudy (anonymous):

The width of a rectangle is 5 feet, and the diagonal is 8 feet. Which is the area of the rectangle? (Round to nearest hundredth.) A)14.14 square feet B)20 square feet C)31.22 square feet C)40 square feet

OpenStudy (anonymous):

i can help u

OpenStudy (anonymous):

@caliallmon

OpenStudy (anonymous):

okay. Please do

OpenStudy (anonymous):

area of triangle = 0.5 * 8 * 5 = 20 square feet therefore area of rectangle= 2*20= 40 square feet!

OpenStudy (anonymous):

Otayy, thank you very much.

rishavraj (rishavraj):

its like